OpenStudy (anonymous):

Simplify: the square root of 2 (x − the square root of 5)

thomaster (thomaster):

so \(\large 2^2(x-5^2)\) ? what is \( 2^2\)? and \(5^2\)?

thomaster (thomaster):

oh i'm sorry it's square root yea

thomaster (thomaster):

\(\Large\sqrt2(x-\sqrt5)\) You could get rid of the parenthesis. \(\Large\sqrt2*x\) and \(\Large\sqrt2*-\sqrt5)\)

OpenStudy (anonymous):

2 times x times the square root of 2 10 the square root of 2 times x the square root of 2 times x − the square root of 10 x times the square root of 2 − the square root of 10

OpenStudy (anonymous):

these are the answers they give me

OpenStudy (anonymous):

would it be C?

thomaster (thomaster):

I think it's pretty obvious that \(\large\sqrt2*2=\sqrt2x\) But do you know what \(\large\sqrt2*-\sqrt5\) is?

thomaster (thomaster):

sorry \(\large\sqrt2*x=\sqrt2x\)

OpenStudy (anonymous):

sqrt -10?

OpenStudy (anonymous):

-sqrt(10) i mean

thomaster (thomaster):

Yes It will be \(\large -\sqrt{10}\)

thomaster (thomaster):

so now we have \(\Large \sqrt2x-\sqrt{10}\)

OpenStudy (anonymous):

Is that it?

thomaster (thomaster):

So both C and D are right, since it doesn't matter when you switch a multiplication. the square root of 2 times x is the same as x times the square root of 2 But i would go for C :)

OpenStudy (anonymous):

OKay thank you! (:

thomaster (thomaster):

No problem, sorry for the many typos :P